I dont know if it has to do much with it but ok.
find the length s and area A.Round answers to three decimal places. In a circle.
r=2 ft
Angle= Pi/3

Answer by Alan3354(69443)

I dont know if it has to do much with it but ok.
find the length s and area A.Round answers to three decimal places. In a circle.
r=2 ft
Angle= Pi/3
------------------
If s is the arc length:
s = r*angle = 2pi/3 feet
------
Area = pi*r^2*(pi/3)/2pi
= pi*r^2/3